OpenStudy (anonymous):

Help! Write 5^0 X 5^-12 using positive exponents.

OpenStudy (anonymous):

\[5^{0} \times 5^{-12}\]

Nnesha (nnesha):

exponent rule when we multiply same bases we should `add` their exponents \[\huge\rm x^m \times x^n=x^{m+n}\] and when we divide same bases we should `subtract` their exponents \[\huge\rm \frac{ x^m }{ x^n }=x^{m-n}\]

OpenStudy (anonymous):

So would it be \[5^{-12}\] ? Cause i need a positive exponent

OpenStudy (anonymous):

or use positive exponents

Nnesha (nnesha):

yes right next step \[\huge\rm x^{-m}=\frac{ 1 }{ x^m }\]

Nnesha (nnesha):

you just to know those exponent rules that's it :=)

OpenStudy (anonymous):

okay so the final would be \[\left(\begin{matrix}1 \\ 5^{12}\end{matrix}\right)\] ?

Nnesha (nnesha):

yes right!

OpenStudy (anonymous):

Okay thanks for the third time lol :)

Nnesha (nnesha):

haha yw! :=)